OpenStudy (anonymous):

how do I solve g(x)=5cos^2((pi)t) ??

OpenStudy (anonymous):

derivative

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

im sorry but I don't understand u, loser66...

OpenStudy (anonymous):

yes sorry, sorry, sorry :(

OpenStudy (anonymous):

its g(t)=5cos^2((pi)t)

OpenStudy (anonymous):

kinda...

OpenStudy (anonymous):

u can show me ur way, itll probably be helpful...

OpenStudy (anonymous):

ok, I think I see it...

OpenStudy (anonymous):

hold on a lil im trying to understand u a lil more

OpenStudy (anonymous):

what happened to the 5?

OpenStudy (anonymous):

lol its ok, but im actually really lost :((

OpenStudy (anonymous):

2

OpenStudy (anonymous):

sorry I forgot the x, I get that..

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

7x^6

OpenStudy (anonymous):

sorry and its 2cos ??

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

-sin

OpenStudy (anonymous):

so -2sin?

OpenStudy (anonymous):

so -2cos(t)sin(t) ??

OpenStudy (anonymous):

2 ?

OpenStudy (anonymous):

2cos ??

OpenStudy (anonymous):

oh, so -2sin(t) ?

OpenStudy (anonymous):

ok I get that :)

OpenStudy (anonymous):

ummm... -4sin(t) ?

OpenStudy (anonymous):

ok :(

OpenStudy (anonymous):

\[g(t)=5\cos^2(\pi(t))\]

OpenStudy (loser66):

wait a minute,

OpenStudy (anonymous):

ok

OpenStudy (loser66):

I am so sorry. i think this way is easier than chain rule, but it seems I confuse you. If you don't get it, let tag other one. @zepdrix

zepdrix (zepdrix):

what, what's going on? :U

OpenStudy (loser66):

I tried to help but my dummy way confuse the asker, help me friend

OpenStudy (anonymous):

sorry :(

zepdrix (zepdrix):

\[\Large g(x)=5\cos^2\left(\pi t\right)\]Find g'(x)? :O hmm

OpenStudy (anonymous):

no sorry its g(t)

zepdrix (zepdrix):

Ah yes :) my mistake

OpenStudy (anonymous):

sorry, wrote it wrong :/

zepdrix (zepdrix):

So we have to apply the `chain rule` several times. Do you understand that the `outermost function` is the ( )^2? If not, we can write our function like this,\[\Large g(t)=5\left[\cos(\pi t)\right]^2\]

OpenStudy (anonymous):

ok, can we do it like that?

zepdrix (zepdrix):

That power placement on trig functions can be a little tricky :U So when we take our derivative, we start by applying the `power rule` to the `outermost function`.\[\Large g'(t)=5\cdot2\left[\cos(\pi t)\right]\]But the chain rule tells us to multiply by the derivative of the inner function,\[\Large g'(t)=5\cdot2\left[\cos(\pi t)\right]\color{royalblue}{\left[\cos(\pi t)\right]'}\]

OpenStudy (anonymous):

thank you though @Loserr66

OpenStudy (anonymous):

@loser66 ^

zepdrix (zepdrix):

Winner66 \:D/

OpenStudy (loser66):

watever,

OpenStudy (anonymous):

:/

zepdrix (zepdrix):

So we need to take the derivative of that blue portion still. :O Understand the process kinda? It takes a little getting used to.

OpenStudy (loser66):

@zepdrix still have (pi t)'

zepdrix (zepdrix):

Yah we haven't taken the derivative of the cosine part yet :3 (pi t)' coming soon :D

OpenStudy (anonymous):

derivative of cos(pi t)= -sin(pi) ??

zepdrix (zepdrix):

Whenever we take the derivative of our `outer function`, we leave the inside alone. So in this case, cosine is our outer function. Differentiating cosine gives us -sine, yes. But the inside should stay the same.

zepdrix (zepdrix):

\[\Large g'(t)=5\cdot2\left[\cos(\pi t)\right]\color{orangered}{\left[-\sin(\pi t)\right]}\]Ok good! But we have to apply the chain rule again! :O\[\Large g'(t)=5\cdot2\left[\cos(\pi t)\right]\color{orangered}{\left[-\sin(\pi t)\right]}\color{royalblue}{\left(\pi t\right)'}\]

OpenStudy (anonymous):

si its, 10(cos(pi t))(-sin(pi t))(pi) ??

zepdrix (zepdrix):

Yes, good. Now since multiplication is commutative, we can rearrange things so it looks a little nicer. Bring the pi and negative sign to the front.

OpenStudy (anonymous):

\[-10\pi(\cos(\pi t))(-\sin(\pi t))\]

OpenStudy (anonymous):

?? yes, orr no?:O

zepdrix (zepdrix):

Woops, when you moved the negative sign to the front, you accidentally left it on the sine also lol. \[\large g'(t)=-10\pi \cos(\pi t)\sin(\pi t)\]

OpenStudy (anonymous):

oh true, sorry about that :o

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

and that's it?? :O -_____-

OpenStudy (loser66):

thaaaatttt 's iiiiiit??? nope, I am struggling with it, not just "that'/s/it" friend

OpenStudy (anonymous):

lol what??

OpenStudy (loser66):

ok, I am mean person!! sorry, I take it back, yeah, you are right, "That's it" thanks @zepdrix

zepdrix (zepdrix):

:x

OpenStudy (anonymous):

so @zepdrix dats da final answer?

zepdrix (zepdrix):

ya :o

OpenStudy (anonymous):

I feel sooo stupid!!! :(

OpenStudy (loser66):

ya dat wud da final answer hahahaha.... how fun it is, teen words, right?

OpenStudy (anonymous):

haha sorry :p just used to them...

OpenStudy (anonymous):

thank you both! @zepdrix @loser66 :)

zepdrix (zepdrix):

water you talkking about loser? o3o doesn't everione talk like dat?